[figure: 9.7 cm]

Dielectric breakdown strength in magnetic fluids

P. Kopčanský^{a}, L. Tomčo^{b}, K. Marton^{c}, M. Koneracká^{a}, I. Potočová^{a},

M. Timko^{a}, J. Jadżyn^{d} and G. Czechowski^{d}

aInstitute of Experimental Physics of Slovak Academy of Sciences, Košice, Slovakia

bAir Force Academy of gen. M. R. Štefánik, Košice, Slovakia

cTechnical University, Faculty of Electrical Eng. and Informatics, Department of High

Voltage Technology, Košice, Slovakia dInstitute of Molecular Physics, Polish Academy of Sciences, Poznaň, Poland

The application of external magnetic field on the magnetic fluids induces needle like structuralization of magnetic particles. The influence of this aggregation of magnetic particles in an external magnetic field \overrightarrow{H} on the magneto-dielectric behaviour and dielectric breakdown strength of a transformer oil based magnetic fluid was measured. Magneto-dielectric behaviour showed anisotropy factor g = 3.4 ( g is defined as ratio of change of dielectric permittivity for \overrightarrow{H}∥\overrightarrow{E} and for \overrightarrow{E}⊥\overrightarrow{H} ) in contrary with theoretical prediction g = 2. This fact indicates developed induced structuralization of magnetic particles. The observed decrease of dielectric breakdown strength ( for H = 0 and \overrightarrow{H}⊥\overrightarrow{E}, where \overrightarrow{E} is applied electric field intensity ) vs distance between electrodes is found to be in agreement with a theory of semi-classical bond percolation for a nonlinear composite material. The oscillation of dielectric breakdown strength vs distance between electrodes was observed when \overrightarrow{E}∥\overrightarrow{H}, i.e. needle-like aggregates are oriented in direction of the applied electric field intensity. This effect is associated with size phenomena probably.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : P. Kopčanský

Address for correspondence : Institute of Experimental Physics of Slovak Academy of Sciences Watsonova 47 043 53 Košice Slovak republic

Email address : kopcan@saske.sk